Find the general solution of the given system of equations: 1 1 6 1 6 1 6 1 1 1
Answer #1

we are given matrix as

A=\begin{bmatrix} 1 &1 &6 \\ 1& 6 &1 \\ 6& 1 &1 \end{bmatrix}

Firstly, we will find eigenvalues

\det \left(\begin{pmatrix}1&1&6\\ 1&6&1\\ 6&1&1\end{pmatrix}-\lambda\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\right)= -\lambda^3+8\lambda^2+25\lambda-200

-X3 +812 + 25x – 200 = 0

\lambda=8,\:\lambda=-5,\:\lambda=5

For T=8:

v_1=\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}

For T=-5:

v_2=\begin{pmatrix}-1\\ 0\\ 1\end{pmatrix}

For T=5:

v_3=\begin{pmatrix}1\\ -2\\ 1\end{pmatrix}

now, we can find solution

x(t)=C_1e^{8t}\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}+C_2e^{-5t}\:\begin{pmatrix}-1\\ 0\\ 1\end{pmatrix}+C_3e^{5t}\:\begin{pmatrix}1\\ -2\\ 1\end{pmatrix}.............Answer